And ancient Greek man it was said to have lived 1/4 of his life as a boy, 1/5 as a youth, 1/3 as a man, and spent the last 13 years as an old man. How old was he when he died?
Answer by jorel555(1290)

The various stages of his life were 1/4, 1/5, and 1/3. Together, they make:
15/60+12/60+20/60=47/60
The remaining years of his life were, therefore, 1-(47/60), or 13/60ths of his whole life. So:
13x/60=13
x=60
The man lived 60 years. ☺☺☺☺